OpenStudy (anonymous):

An inlet pipe can fill a large water tank in 123 min. An outlet pipe can empty the same water tank in 248 minutes. Given an empty water at 11 am on Tuesday, activate both pipes. what time on the clock will it be before the tank is filled? am or pm and day of week

OpenStudy (anonymous):

It's filled after x minutes: \[ (\frac{1}{123} - \frac{1}{248}) x = 1 \] so \[x = \frac{30504}{125} \] minutes later. Then ( http://www.wolframalpha.com/input/?i=30504%2F125+minutes+after+11AM) the tank will be filled at 3:04:01 PM on Tuesday.